OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (cwrw238):

geometric series is a1, a1*r , a1*r^2 .... so a1r^2 / a1 = a3/a1 = r^2 so r^2 = 25 / 625 = 1/25 6th term = a1* r^5 = 625 * (1/25)^5 ur calculator will help you work this out